Tuesday, Mar. 08: DAY Shift
From Hall A Wiki
Go back to Daily Run Plan Daily Run Plans
- Production data (Kin48_3)
- Elastic measurement during DAY
- 6:30 AM: Beam Off, Call MCC to put Hall A in Controlled Access Starting Ramping HRS-L magnets to 0
- 7:30 AM:Controlled Access: Tech in Hall, Change HRS-L Polarity, Move Angle to 40.0 deg, Set Momentum = 2.5106 GeV/c, Q1 = 1900 Amp
- 8:30 AM: Elastic Angle Calibration: DVCS Trigger s0+s2m only in HRS trigger, Elastics with Calo in current Kin_048_3 setting [here], 15 microAmp
- 9:30 AM Escorted Access: Move Calo to 6 m, 14.0 deg - 10:00 AM Elastic Energy Calibration data 20 microAmp
- 10:30 AM Escorted Access: Move Calo to 13.1 deg
- 11:00 AM Elastic Energy Calibration data
- 11:30 AM Escorted Access: Move Calo to 12.25 deg
- 12:00 PM Elastic Energy Calibration data
- 12:30 PM Escorted Access: Move Calo to Kin_48_2 setting: 15.184 deg, D=2.0 m, 10 micro Amp
- 01:00 PM Angle Calibration Data (This puts the central elastic 7.1 GeV electrons 2.0 blocks from the physical edge of the calorimeter.)
- 02:00 PM Controlled Access? Ramp down HRS-L Polarity, Angle change HRS-L to 20.244 deg to p = 3.996 GeV/c, Q1 = 1900 Amp
- 03:00 PM Production DVCS at KIN_48_2[here]
- Back to Production run, Kin48_2
- After Alex's work, Compton laser will be turned OFF !
- Polarized beam at 9 uA on LH2 15 cm!
- DVCS is running kin_48_2 [here]
- LHRS is partially detuned: [here]
- LHRS=3.996GeV, Q1/Q2/Q3=1916.640(readout:1900)/1600/1600 (Q1 regul.OFF)
- PreScaleR : PS1=PS2=PS3=1, && PS8=200
- PreScaleL : PS9 =1, + DVCS config.
- beam locks: no position lock, energy lock in Arc 2: [here]
- DVCS daq configuration is [this one], we are measuring about 2% DT at 15uA.
- After beam lost for considerable time (like beam study, pass change,...)
- Ask MCC to run 5uA wth carbon hole target and make sure you can see the hole through Spot++
Please check more info on the white board in the Counting House ~!!
MCC Beam Team Meeting on Tuesday 1:30PM
- Julie note from Beam Team Meeting at 1:30PM, Mar.8, 2016 at MCC.
- Purpose: Plans to run at 60uA
- Current issues:
- There is no consensus on why the beam is limited at 30 uA.
- 1/ The limitation is in the injector - 2/ In the RF separator
- 1/ Removing/Changing Wien filter (horizontal-> vertical) ?? - 2/ Run with a magnetic separation ??
- Problem foresee:
- 1/ ?? - 2/ Compatible with Hall D taking data
- GMp expert inform:
- 1/ Polarization is not matter - 2/ High current 60 uA should be deliverable for calibration and production both 4pass and 5pass
- Concluded Plan:
- Try 60 uA Thursday morning up to noon